Given the genotype AaBbCcDdEE

a) How many kinds of gametes can this individual form?

b) If selfing occurred, what fraction of the progeny would have the following genotype: AaBBccDdEE?

A cross between two plants produces progeny which, through subsequent breeding experiments, are shown to fall into 27 genotypic classes. Assuming that complete dominance operates at all of the loci involved and that each locus has two alleles,

  1. a) At how many different loci are the parents heterozygous?

  2. b) How many phenotypes would these progeny be expected to show?
Answer #1

1) a)

AaBbCcDdEE

this is heterozygous at 4 loci so number of different kinds of gametes it can produce= 2^ number of heterozygous loci=2^4=16

b)

Aa * Aa

A a
A AA Aa
a Aa aa

fraction of Aa= number of Aa/total number=2/4=1/2

Bb* Bb

B b
B BB Bb
b Bb bb

the fraction of BB= number of BB/total number=1/4

Cc* Cc

C c
C CC Cc
c Cc cc

fraction of cc= number of cc/total number=1/4

Dd * Dd

D d
D DD Dd
d Dd dd

fraction of Dd= number of Dd/total number=2/4=1/2

EE * EE

EE ( all the F1 progenies are EE)

fraction of progeny with AaBBccDdEE genotype= fraction of Aa fraction of BB fraction of cc fraction of Dd fraction of EE= 1/2 1/41/41/21=1/64